# Two-Channel Filter Banks on Joint Time-Vertex Graphs with Oversampled Graph Laplacian Matrix

### v1

@@ VII Conclusion
In this work, we proposed a novel two-channel filter bank for joint time-vertex signal processing. An oversampled graph Laplacian operator was formulated and extended from vertex-domain filter banks to a joint oversampled graph filter bank. Based on this operator, a two-channel oversampled filter bank was designed using a $K$-coloring strategy, enabling the construction of oversampled bipartite graphs that preserve all temporal and spatial edges. Extensive experiments on both signal reconstruction and image denoising validate the effectiveness of the proposed filter banks, demonstrating notable improvements in denoising performance compared to existing methods. Future work will focus on optimized filter design, extensions to directed and dynamic graphs, and broader applications in large-scale time-varying image data.

### v2

@@ VII Conclusion
In this work, we proposed a novel two-channel filter bank for joint time-vertex signal processing. An oversampled graph Laplacian operator was formulated and extended from vertex-domain filter banks to a joint oversampled graph filter bank. Based on this operator, a two-channel oversampled filter bank was designed using a $K$-coloring strategy, enabling the construction of oversampled bipartite graphs that preserve all temporal and spatial edges. Extensive experiments on both signal reconstruction and image denoising validate the effectiveness of the proposed filter banks, demonstrating notable improvements in denoising performance compared to existing methods. Future work will focus on optimized and adaptive filter design, extensions to directed and dynamic graphs, and broader applications in large-scale time-varying signal data.
@@ Appendix B Proof of Theorem 1
Acknowledgments
In the spirit of reproducible research, MATLAB code examples are available at https://github.com/Zhangyubit/JOGLM.